# On the common index jump theorem and further developments

### v1

@@ Abstract
In [LZ02], Long and Zhu established originally the common index jump theorem (CIJT) for symplectic paths in 2002, which was later generalized to its enhanced version (ECIJT) by Duan, Long and Wang in [DLW16] in 2016. Started from [GGM18] of 2018 and [GG20] of 2020, and finally in [CGG24] of 2024, a similar index theorem was obtained, i.e., Theorem 3.3 of [CGG24], which was called the index recurrence theorem there. In this short note, we give detailed proofs to show that the first 4 assertions dealing with index iterations in the total of 5 assertions in Theorem 3.3 of [CGG24] actually coincide completely with results in (ECIJT) of [DLW16].
@@ 1 Introduction
In the paper [LZ02] of Long and Zhu published in 2002, the common index jump theorem (CIJT for short below) was established, and since then this theorem has become a powerful tool in the study of many problems related to periodic solutions of Hamiltonian systems.
@@
Note that in the paper [DLW16] of Duan, Long and Wang published in “Calculus of Variations” in 2016, CIJT was improved to the enhanced common index jump theorem (ECIJT for short below), i.e., Theorem 3.5 of [DLW16], with some more precise formula on index iterations. Then in recent [DLLW24] of Duan, Liu, Long and Wang in 2024, ECIJT was further generalized.
@@
On the other hand, In [GGM18] of Ginzburg, Gürel and Macarini published in 2018, a so-called index recurrence theorem, Theorem 4.1 there, was obtained, and the authors wrote specially that “This result can be deduced from the so-called enhanced common index jump theorem due to Duan, Long and Wang [DLW16] (see also [Lo02], [LZ02])”. In [GG20] of Ginzburg and Gürel published in 2020, its Theorem 5.2 which was also called the index recurrence theorem, was obtained. In [CGG24] of Cineli, Ginzburg and Gürel appeared as arXiv:2410.13093v1 on October 16, 2024 and its revision arXiv:2410.13093v2 on July. 17, 2025, its Theorem 3.3 was obtained and was also called index recurrence theorem too, whose proof was a replenished version of and was more rigorous than that of Theorem 5.2 in [GG20]. Such a theorem became one of the main tools in the proof of the main multiplicity result on closed Reeb flow orbits in [CGG24].
@@
Note that the Theorem 3.3 in [CGG24] contains total of 5 assertions (IR1)-(IR5), and it needs to be pointed out that the first 4 assertions there coincide completely with the identities (3.20)-(3.22) of Theorem 3.5 (ECIJT) and estimates (3.42) and (3.43) in Remark 3.6 in [DLW16]. Note also that the proofs of (IR1)-(IR4) in [CGG24] occupied the major part in the proofs of Theorem 3.3 in [CGG24]. On the other hand because all the assertions in Theorem 5.2 in [GG20] are covered by these (IR1)-(IR4) in [CGG24], and thus assertions of Theorem 5.2 in [GG20] coincide also with those proved in [DLW16] published in 2016 already too.
@@
Because in [CGG24], the authors used the Arnold-Givental normal forms of real quadratic forms established in [AG], and in [DLW16] the authors used the normal forms and basic normal forms of real symplectic matrices established in [LD00], [LH99] and [Lo02] (cf. Subsections 1.4-1.9 on pp.16-47 and Chapter 8 in [Lo02] for details), the coincidences between above mentioned (IR1)-(IR4) of Theorem 3.3 in [CGG24] and results already proved in [DLW16] are not that obvious for usual readers. This short paper is devoted to give detailed proofs for such coincidences, which are contained specially in the 4 Propositions below.
@@
In the following, when the page numbers of [CGG24] appear we are using those in arXiv:2410.13093v2 on July. 17, 2025.
2 Theorem 3.3 and Remark 3.6 of [DLW16] as well as Theorem 3.3 of [CGG24]
Note that in Theorem 3.5 on p.145 of [DLW16], the authors proved the following theorem,
@@
Theorem 3.5 in [DLW16]. (The enhanced common index jump theorem for symplectic paths) Let ${\gamma}_{k}\in{\cal P}_{\tau_{k}}(2n)$ for $k=1,\ldots,q$ be a finite collection of symplectic paths. Let $M_{k}={\gamma}_{k}(\tau_{k})$. We extend ${\gamma}_{k}$ to $0,+\infty)$ by (3.9) of [DLW16] inductively. Suppose
$\hat{i}({\gamma}_{k},1)>0,\qquad\forall\;k=1,\ldots,q.\hskip 227.62204pt[DLW16,(3.17)].$
@@
$\check{m}\equiv\check{m}({\gamma}_{1},\ldots,{\gamma}_{q})=\min\{\check{M}_{k}\;|\;1\leq k\leq q\}.\hskip 170.71652pt[DLW16,(3.18)].$
@@
$\nu({\gamma}_{k},2m_{k}-m)=\nu({\gamma}_{k},2m_{k}+m)=\nu({\gamma}_{k},1),\;\forall\;1\leq k\leq q,\;1\leq m<\check{m},\hskip 28.45274pt[DLW16,(3.19)],$
@@
$\nu({\gamma}_{k},2m_{k}-m)=\nu({\gamma}_{k},2m_{k}+m)=\nu({\gamma}_{k},m),\hskip 128.0374pt[DLW16,(3.20)],$
$i({\gamma}_{k},2m_{k}+m)=2N+i({\gamma}_{k},m),\hskip 187.78836pt[DLW16,(3.21)],$
$i({\gamma}_{k},2m_{k}-m)=2N-i({\gamma}_{k},m)-2(S_{M_{k}}^{+}(1)+Q_{k}(m)),\hskip 85.35826pt[DLW16,(3.22)],$
$i({\gamma}_{k},2m_{k})=2N-(S_{M_{k}}^{+}(1)+C(M_{k})-2{\Delta}_{k}),\hskip 128.0374pt[DLW16,(3.23)],$
where as in [LZ02], we let
${\Delta}_{k}=\sum_{0<m_{k}{\theta}/\pi\}<{\delta}}S_{M_{k}}^{-}(e^{\sqrt{-1}{\theta}}),\hskip 213.39566pt[DLW16,(3.24)],$
@@
$Q_{k}(m)=\sum_{{\theta}\in(0,2\pi),\;e^{\sqrt{-1}{\theta}}\in{\sigma}(M_{k})\atop\{\frac{m_{k}{\theta}}{\pi}\}=\{\frac{m{\theta}}{2\pi}\}=0}S_{M_{k}}^{-}(e^{\sqrt{-1}{\theta}}).\qquad\hskip 142.26378pt[DLW16,(3.25)]$
Note that in (3.42) and (3.43) contained in Remark 3.6 on p.145 in [DLW16], the authors proved also the following two estimates.
Remark 3.6 (ii) in [DLW16]. By (4.10) and (4.40) in [LZ02] (cf. (11.1.10) and (11.2.14) of [Lo02]), we have
$m_{k}=\left(\left[\frac{N}{\bar{M}\hat{i}({\gamma}_{k},1)}\right]+\chi_{k}\right)\bar{M},\qquad 1\leq k\leq q,\hskip 85.35826pt[DLW16,(3.42)]$
where $\chi_{k}=0$ or $1$ for $1\leq k\leq q$, $\bar{M}$ is a positive integer such that $\frac{\bar{M}{\theta}}{\pi}\in{\bf Z}$ whenever $e^{\sqrt{-1}{\theta}}\in{\sigma}(M_{k})$ and $\frac{{\theta}}{\pi}\in{\bf Q}$ for some $1\leq k\leq q$, and we set $\bar{M}=1$ if no such eigenvalues exist. Furthermore, by (4.20) in Theorem 4.1 of [LZ02] (cf.(11.1.20) of [Lo02]), for any ${\epsilon}>0$, we can choose $N$ and $\{\chi_{k}\}_{1\leq k\leq q}$ such that
$\left|\left\{\frac{N}{\bar{M}\hat{i}({\gamma}_{k},1)}\right\}-\chi_{k}\right|<{\epsilon},\qquad 1\leq k\leq q.\hskip 113.81102pt[DLW16,(3.43)]$
@@
On the other hand, in p.30 of [CGG24] put into arXiv on Oct. 16, 2024 and then on July.17,2025, under similar conditions the authors proved their Theorem 3.3 which contains five assertions (IR1)-(IR5). Here we are interested in the following first four assertions (IR1)-(IR4) on the index iteration theories of symplectic matrix paths started from the identity matrix.
Theorem 3.3 in [CGG24]. For any $\eta>0$ and any $\ell_{0}\in{\bf N}$, there exists a sequence $C_{l}\to\infty$ of real numbers and integer sequences $d_{il}\to\infty$ and $k_{ijl}\to\infty$ (as $l\to\infty$) such that for all $i$, $j$ and $l$, and all $\ell\in{\bf Z}$ in the range $1\leq|\ell|\leq\ell_{0}$, we have
@@
Note that all of (IR1)-(IR4) are in fact claims on only one single symplectic path, and clearly (IR4) is a simple consequence of (IR2)-(IR3). Thus the authors of [CGG24] pointed out specially that their proofs for (IR1)-(IR3) can be simplified to the following three properties on a single symplectic path.
@@
Next we explain the details of the comparisons of results contained in these 2 papers related to (IR1)-(IR4) in the following 4 sections.
3 The coincidence of ${\beta}_{-}({\gamma})$ in [CGG24] and $S_{{\gamma}(1)}^{-}(1)$ in [Lo02]
To get precise definitions of certain quantities used in [CGG24] and understand the proof of Theorem 3.3 in [CGG24], we make first the following concrete computations, and then prove the Proposition 1 below.
As in p.27 of [CGG24], using the normal form of symplectic matrices in Section 2.4 of [AG], the authors considered a totally degenerate matrix $A\in{\rm Sp}(2m)$. Then $A$ can be written into $A=\exp(JQ)$, where all eigenvalues of $JQ$ are zero. The quadratic form $Q$ can be symplectically decomposed into a sum of terms of the following three types according to Section 2.4 of [AG]:
@@ 3 The coincidence of ${\beta}_{-}({\gamma})$ in [CGG24] and $S_{{\gamma}(1)}^{-}(1)$ in [Lo02]
where $\nu(A)$ is the geometric multiplicity of the eigenvalue $1$ of $A$ as usual (cf. Definition 5.1.1 on p.111 of [Lo02]).
As shown below in our discussion, such a quadratic form $Q$ corresponds to a symplectic path $\Phi\in C([0,1],{\rm Sp}(2n))$. Then as described in p.27 of [CGG24], the authors decompose $\Phi$ into a sum of two paths $\Phi=\Phi_{0}\oplus\Psi$ such that $\Phi_{0}(1)\in{\rm Sp}(2n_{0})$ is totally degenerate and $\Psi(1)\in{\rm Sp}(2n_{1})$ is non-degenerate. In particular $n_{0}$ is the algebraic multiplicity of the eigenvalue $1$ of $\Phi(1)$ and $n_{0}+n_{1}=n$. Then the authors of [CGG24] defined first
@@
Note that the values of the function $b_{0}$ introduced in p.27 of [CGG24] is defined by $b_{0}(Q)=\nu_{0}(Q_{0})$. From the definition of $Q_{0}$ in the above (ii), we obtain
@@
Here and below, according to Examples 2) in [AG,p.5], the standard symplectic matrix on the coordinate space ${\bf R}^{2n}$ is given by $J=\left(\matrix{0&-I_{n}\cr I_{n}&0\cr}\right)$ where $I_{n}$ is the $n\times n$ identity matrix on ${\bf R}^{n}$. According to the Corollary in p.5 of [AG], the Darboux coordinates is written as $(p_{1},\ldots,p_{d},q_{1},\ldots,q_{d})\in{\bf R}^{2n}$. Note that these choices coincides with those chosen in [Lo02] which are related to the orientation in the discussions here, and are important in the computations blow.
@@
By the study in Section 1.4 on pp.16-24 and Theorem 1.8.10 on p.41 of [Lo02], the matrix ${\gamma}(1)$ can be connected to a decomposition of basic normal forms in ${\Omega}_{1}^{0}({\gamma}(1))$, where the homotopy set ${\Omega}_{1}(M)$ respect to the eigenvalue $1$ is defined by
@@
and ${\Omega}_{1}^{0}(M)$ is the path connected component of ${\Omega}_{1}(M)$ containing $M$ defined in [Lo99b] and can be find in Definition 1.8.5 on p.38 of [Lo02]. Specially the above matrix ${\gamma}(t)$ with $t\in[0,1]$ is in the form of the matrix $N_{d}(1,0)$ defined in (1.4.1)-(1.4.5) with $b=0$ on pp.17-18 in [Lo02] which is written as
@@
Thus by Theorem 1.4.1 on p.18 of [Lo02], we obtain
@@
Using the basic normal forms (defined in [Lo99a], [Lo00]) as in (1.8.15)-(1.8.17) on p.41 in [Lo02], by Theorem 1.8.10 on p.41 of [Lo02] again, it yields
@@
In this case, authors of [CGG24] defined $b_{0}(Q)$ to be the number of $Q_{0}$ in the decomposition of $Q$. When $Q=Q_{0}$ given by (3.1) they defined
@@
Using the splitting numbers $S^{\pm}_{M}({\lambda})$ defined in [Lo99b] (cf. Definition 9.1.4 on p.191 of [Lo02]), by the List 12 on p.198 of [Lo02] we have
@@
In this subcase, by the definition in [AG] and [CGG24], the quadratic form $Q_{\pm}$ is defined on ${\bf R}^{2}$ as $Q_{{\epsilon}}$ with ${\epsilon}=1$ or 1-1 and the coordinates on ${\bf R}^{2}$ is given by $(p_{d},q_{d})$. In this case, from
@@
with $N_{1}(1,-{\epsilon})=\left(\matrix{1&-{\epsilon}\cr 0&1\cr}\right)$. Therefore by List 9.1.12 on p.198 of [Lo02], we obtain
@@
In this subcase, for ${\epsilon}=\pm 1$, by the definition in [AG] and [CGG24] the quadratic form $Q_{{\epsilon}}$ is given by $Q_{{\epsilon}}={\epsilon}(Q_{0}+p_{d}^{2}/2)$. Let ${\gamma}_{{\epsilon}}(t)$ be the fundamental solution of the initial value problem of the linear Hamiltonian system (3.16)-(3.17).
Then considering the above study on $Q_{0}$ on ${\bf R}^{2d}$ and $Q_{\pm}$ on ${\bf R}^{2}$, specially (3.3) and (3.10), by the study in Sections 1.8 and 1.9 on pp.36-47 in [Lo02], the basic normal form decomposition of ${\gamma}_{\pm 1}$ can be written as follows,
@@
Because the invariants in (3.3) and the splitting numbers $S_{M}^{\pm}({\lambda})$ in [Lo99b] and Definition 9.1.4 on p.191 of [Lo02] are symplectic additive, by (3.25) in the subcase 1.3.2 we obtain
@@
4 The (IR1) in Theorem 3.3 of [CGG24] is a direct consequence of Remark 3.6 of [DLW16]
In Section 3 of [CGG24], as usual the authors denote the set of all non-degenerate symplectic matrices in ${\rm Sp}(2m)$ with $m\in{\bf N}$ by ${\rm Sp}^{\ast}(2m)$. Then for a symplectic path $\Phi$ with end point $\Phi(1)\in{\rm Sp}^{\ast}(2m)$, the end point matrix $\Phi(1)$ can be connected to I-I in ${\rm Sp}^{\ast}(2m)$ by a path $\Psi$ lying entirely in ${\rm Sp}^{\ast}(2m)$. Let $\Phi^{\prime}$ be the concatenating path of $\Phi$ and $\Psi$. Then as usual the Conley-Zehnder index of $\Phi$ is defined by
@@
Then the authors of [CGG24] defined two indices for $\Phi\in\widetilde{{\rm Sp}}(2m)$ by
@@
Our next result shows that the relations of the property (IR1”) with the indices defined in [Lo02] as well as [DLW16] are as follows,
Proposition 2. (i) The assertion (IR1”-1) in (IR1) are direct consequence of (3.42)-(3.43) in [DLW16].
(ii) The assertion (IR1”-2) in (IR1) was already proved in (1) of Theorem 1.1 of [LL00] (cf. $1^{o}$ of Theorem 10.1.2 on p.213 of [Lo02]).
Proof. Note first that using definitions of $\mu_{\pm}(\Phi)$ in (4.2), by Corollary 6.1.12 and Definition 6.1.13 on p.144 of [Lo02], it yields immediately

### v2

@@ Abstract
In [18] published in “Annals of Math.”, Long and Zhu established originally the common index jump theorem (CIJT) for symplectic paths in 2002, which has played an important role in later studies on periodic solution orbits for Hamiltonian systems, Reeb flows, and geodesic problems. This (CIJT) was generalized to its enhanced version (ECIJT) by Duan, Long and Wang in [3] in 2016. Started from [5] of 2020, and finally in [2] of 2024, a similar index theorem was obtained, i.e., Theorem 3.3 of [2], which was given the name “index recurrence theorem” there. In this short note, we give detailed proofs to show that the major assertions, i.e., the first 4 assertions in the total of 5 assertions, in Theorem 3.3 of [2] as well as all the assertions in [5] actually coincide completely with results in (ECIJT) of [3].
Keywords: Common index jump theorem, enhanced version, comparison, iterated index computation, estimates.
2020 Mathematics Subject Classification: 58E05, 34C25.
Abbreviated title: Common index jump theorem and its extensions.
@@ 1 Introduction
In the paper [18] of Long and Zhu published in 2002, the common index jump theorem (CIJT for short below) was established, and since then this theorem has become a powerful tool in the study of many problems related to periodic solutions of Hamiltonian systems.
@@
Note that in the paper [3] of Duan, Long and Wang published in “Calculus of Variations” in 2016, CIJT was improved to the enhanced common index jump theorem (ECIJT for short below), i.e., Theorem 3.5 of [3], with some more precise formula on index iterations. Then in recent [4] of Duan, Liu, Long and Wang in 2024, ECIJT was further generalized.
On the other hand, In [6] of Ginzburg, Gürel and Macarini published in 2018, a so-called index recurrence theorem, i.e., Theorem 4.1 there, was obtained, and the authors wrote in their abstract of [6] specially that “On the combinatorial side of the question, we revisit and reprove the enhanced common jump theorem (i.e., the (ECIJT) Theorem 3.5) from [3] and interpret it as an index recurrence result (i.e., the Theorem 4.1 in [6])”.
In [5] of Ginzburg and Gürel published in 2020, its Theorem 5.2 was obtained, which was also called the index recurrence theorem. Note that although the Theorem 4.1 of [6] assumed the strongly non-degenerate condition on the symplectic paths studied, but the Theorem 3.5 and Remark 3.6 in [3] published in 2016 used no any non-degeneracy conditions and allow degenerate symplectic paths, which thus covered already completely all the assertions in Theorem 5.2 of [5] published in 2020, which is 4 years later than [3].
@@
In [2] of Cineli, Ginzburg and Gürel appeared as arXiv:2410.13093v3 on Oct. 11, 2025, the Theorem 5.2 of [5] was extended to the Theorem 3.3 of [2] and was also called index recurrence theorem too, whose proof was a replenished version of and was more rigorous than that of Theorem 5.2 in [5]. This Theorem 3.3 became one of the main tools in the proof of the main multiplicity result on closed Reeb flow orbits in [2].
@@
Note that the Theorem 3.3 in [2] contains 5 assertions (IR1)-(IR5) in total, and it needs to be pointed out that the first 4 assertions there coincide completely with the identities (3.20)-(3.22) of Theorem 3.5 (ECIJT) and estimates (3.42) and (3.43) in Remark 3.6 in [3] already published in 2016. Note also that the proof of (IR1)-(IR4) in [2] occupied the major part in the proofs of Theorem 3.3 in [2]. On the other hand because all the assertions in Theorem 5.2 in [5] are covered by these (IR1)-(IR4) in [2], as mentioned above all the assertions of Theorem 5.2 in [5] coincide also with those proved in [3] published in 2016 already too.
@@
Because for notations and proofs, in [2] the authors used the Arnold-Givental normal forms of real quadratic forms established by Arnold and Givental in [1] of 2001, and in [3] the authors used the normal forms and basic normal forms of real symplectic matrices established by Dong, Han and Long in [16], [17] and [14] (cf. Subsections 1.4-1.9 on pp.16-47 and Chapter 8 in [15] for details), the coincidences between above mentioned (IR1)-(IR4) of Theorem 3.3 in [2] and results already proved in [3] are not that obvious for usual readers. This short paper is devoted to give detailed proofs for such coincidences, which are contained specially in the 4 Propositions below.
@@
In the following, when the page numbers of [2] appear we are using those in arXiv:2410.13093v3 on Oct. 11, 2025.
2 Theorem 3.3 and Remark 3.6 of [3] as well as Theorem 3.3 of [2]
Note that in Theorem 3.5 on p.145 of [3], the authors proved the following theorem,
@@
Theorem 3.5 in [3]. (The enhanced common index jump theorem for symplectic paths) Let ${\gamma}_{k}\in{\cal P}_{\tau_{k}}(2n)$ for $k=1,\ldots,q$ be a finite collection of symplectic paths. Let $M_{k}={\gamma}_{k}(\tau_{k})$. We extend ${\gamma}_{k}$ to $0,+\infty)$ by (3.9) of [3] inductively. Suppose
$\hat{i}({\gamma}_{k},1)>0,\qquad\forall\;k=1,\ldots,q.\hskip 227.62204pt{\rm(3.17)\;of\;[3]}.$
@@
$\check{m}\equiv\check{m}({\gamma}_{1},\ldots,{\gamma}_{q})=\min\{\check{M}_{k}\;|\;1\leq k\leq q\}.\hskip 170.71652pt{\rm(3.18)\;of\;[3]}.$
@@
$\nu({\gamma}_{k},2m_{k}-m)=\nu({\gamma}_{k},2m_{k}+m)=\nu({\gamma}_{k},1),\;\forall\;1\leq k\leq q,\;1\leq m<\check{m},\hskip 28.45274pt{\rm(3.19)\;of\;[3]},$
@@
$\nu({\gamma}_{k},2m_{k}-m)=\nu({\gamma}_{k},2m_{k}+m)=\nu({\gamma}_{k},m),\hskip 128.0374pt{\rm(3.20)\;of\;[3]},$
$i({\gamma}_{k},2m_{k}+m)=2N+i({\gamma}_{k},m),\hskip 187.78836pt{\rm(3.21)\;of\;[3]},$
$i({\gamma}_{k},2m_{k}-m)=2N-i({\gamma}_{k},m)-2(S_{M_{k}}^{+}(1)+Q_{k}(m)),\hskip 85.35826pt{\rm(3.22)\;of\;[3]},$
$i({\gamma}_{k},2m_{k})=2N-(S_{M_{k}}^{+}(1)+C(M_{k})-2{\Delta}_{k}),\hskip 128.0374pt{\rm(3.23)\;of\;[3]},$
where as in [18], we let
${\Delta}_{k}=\sum_{0<\{m_{k}{\theta}/\pi\}<{\delta}}S_{M_{k}}^{-}(e^{\sqrt{-1}{\theta}}),\hskip 213.39566pt{\rm(3.24)\;of\;[3]},$
@@
$Q_{k}(m)=\sum_{{\theta}\in(0,2\pi),\;e^{\sqrt{-1}{\theta}}\in{\sigma}(M_{k})\atop\{\frac{m_{k}{\theta}}{\pi}\}=\{\frac{m{\theta}}{2\pi}\}=0}S_{M_{k}}^{-}(e^{\sqrt{-1}{\theta}}).\qquad\hskip 142.26378pt{\rm(3.25)\;of\;[3]}$
Note that in (3.42) and (3.43) contained in Remark 3.6 on p.145 in [3], the authors proved also the following two estimates.
Remark 3.6 (ii) in [3]. By (4.10) and (4.40) in [18] (cf. (11.1.10) and (11.2.14) of [15]), we have
$m_{k}=\left(\left[\frac{N}{\bar{M}\hat{i}({\gamma}_{k},1)}\right]+\chi_{k}\right)\bar{M},\qquad 1\leq k\leq q,\hskip 85.35826pt{\rm(3.42)\;of\;[3]}$
where $\chi_{k}=0$ or $1$ for $1\leq k\leq q$, $\bar{M}$ is a positive integer such that $\frac{\bar{M}{\theta}}{\pi}\in{\bf Z}$ whenever $e^{\sqrt{-1}{\theta}}\in{\sigma}(M_{k})$ and $\frac{{\theta}}{\pi}\in{\bf Q}$ for some $1\leq k\leq q$, and we set $\bar{M}=1$ if no such eigenvalues exist. Furthermore, by (4.20) in Theorem 4.1 of [18] (cf.(11.1.20) of [15]), for any ${\epsilon}>0$, we can choose $N$ and $\{\chi_{k}\}_{1\leq k\leq q}$ such that
$\left|\left\{\frac{N}{\bar{M}\hat{i}({\gamma}_{k},1)}\right\}-\chi_{k}\right|<{\epsilon},\qquad 1\leq k\leq q.\hskip 113.81102pt{\rm(3.43)\;of\;[3]}$
@@
On the other hand, in p.31 of [2], under similar conditions the authors proved their Theorem 3.3 which contains five assertions (IR1)-(IR5). Here we are interested in the following first four assertions (IR1)-(IR4) on the index iteration theories of symplectic matrix paths started from the identity matrix.
Theorem 3.3 in [2]. For any $\eta>0$ and any $\ell_{0}\in{\bf N}$, there exists a sequence $C_{l}\to\infty$ of real numbers and integer sequences $d_{il}\to\infty$ and $k_{ijl}\to\infty$ (as $l\to\infty$) such that for all $i$, $j$ and $l$, and all $\ell\in{\bf Z}$ in the range $1\leq|\ell|\leq\ell_{0}$, we have
@@
Note that all of (IR1)-(IR4) are in fact claims on only one single symplectic path, and clearly (IR4) is a simple consequence of (IR2)-(IR3). Thus the authors of [2] pointed out specially that their proofs for (IR1)-(IR3) can be simplified to the following three properties on a single symplectic path.
@@
Next we explain the details of the comparisons of results contained in the papers [2] and [3] related to (IR1)-(IR4) in the following 4 sections.
3 The coincidence of ${\beta}_{-}({\gamma})$ in [2] and $S_{{\gamma}(1)}^{-}(1)$ in Long’s book [15] of 2002
To get precise definitions of certain quantities used in [2] and understand the proof of Theorem 3.3 in [2], we make first the following concrete computations, and then prove the Proposition 1 below.
As in p.28 of [2], using the normal form of symplectic matrices in Section 2.4 of [1], the authors considered a totally degenerate matrix $A\in{\rm Sp}(2m)$. Then $A$ can be written into $A=\exp(JQ)$, where all eigenvalues of $JQ$ are zero. The quadratic form $Q$ can be symplectically decomposed into a sum of terms of the following three types according to Section 2.4 of [1]:
@@
where $\nu(A)$ is the geometric multiplicity of the eigenvalue $1$ of $A$ as usual (cf. Definition 5.1.1 on p.111 of [15]).
As shown below in our discussion, such a quadratic form $Q$ corresponds to a symplectic path $\Phi\in C([0,1],{\rm Sp}(2n))$. Then as described in p.28 of [2], the authors decompose $\Phi$ into a sum of two paths $\Phi=\Phi_{0}\oplus\Psi$ such that $\Phi_{0}(1)\in{\rm Sp}(2n_{0})$ is totally degenerate and $\Psi(1)\in{\rm Sp}(2n_{1})$ is non-degenerate. In particular $n_{0}$ is the algebraic multiplicity of the eigenvalue $1$ of $\Phi(1)$ and $n_{0}+n_{1}=n$. Then the authors of [2] defined first
@@
Note that the values of the function $b_{0}$ introduced in p.28 of [2] is defined by $b_{0}(Q)=\nu_{0}(Q_{0})$. From the definition of $Q_{0}$ in the above (ii), we obtain
@@
Here and below, according to Examples 2) on p.5 of [1], the standard symplectic matrix on the coordinate space ${\bf R}^{2n}$ is given by $J=\left(\matrix{0&-I_{n}\cr I_{n}&0\cr}\right)$ where $I_{n}$ is the $n\times n$ identity matrix on ${\bf R}^{n}$. According to the Corollary in p.5 of [1], the Darboux coordinates is written as $(p_{1},\ldots,p_{d},q_{1},\ldots,q_{d})\in{\bf R}^{2n}$. Note that these choices coincides with those chosen in Long’s book [15] of 2002 which are related to the orientation in the discussions here, and are important in the computations blow.
@@
By the study in Section 1.4 on pp.16-24 and Theorem 1.8.10 on p.41 of Long’s book [15], the matrix ${\gamma}(1)$ can be connected to a decomposition of basic normal forms in ${\Omega}_{1}^{0}({\gamma}(1))$, where the homotopy set ${\Omega}_{1}(M)$ respect to the eigenvalue $1$ is defined by
@@
and ${\Omega}_{1}^{0}(M)$ is the path connected component of ${\Omega}_{1}(M)$ containing $M$ defined in [13] and can be find in Definition 1.8.5 on p.38 of [15]. Specially the above matrix ${\gamma}(t)$ with $t\in[0,1]$ is in the form of the matrix $N_{d}(1,0)$ defined in (1.4.1)-(1.4.5) with $b=0$ on pp.17-18 in [15] which is written as
@@
Thus by Theorem 1.4.1 on p.18 of [15], we obtain
@@
Using the basic normal forms (defined in [12], [14]) as in (1.8.15)-(1.8.17) on p.41 in [15], by Theorem 1.8.10 on p.41 of [15] again, it yields
@@
In this case, authors of [2] defined $b_{0}(Q)$ to be the number of $Q_{0}$ in the decomposition of $Q$. When $Q=Q_{0}$ given by (3.1) they defined
@@
Using the splitting numbers $S^{\pm}_{M}({\lambda})$ defined in [13] (cf. Definition 9.1.4 on p.191 of [15]), by the List 12 on p.198 of [15] we have
@@
In this subcase, by the definition in [1] and [2], the quadratic form $Q_{\pm}$ is defined on ${\bf R}^{2}$ as $Q_{{\epsilon}}$ with ${\epsilon}=1$ or 1-1 and the coordinates on ${\bf R}^{2}$ is given by $(p_{d},q_{d})$. In this case, from
@@
with $N_{1}(1,-{\epsilon})=\left(\matrix{1&-{\epsilon}\cr 0&1\cr}\right)$. Therefore by List 9.1.12 on p.198 of [15], we obtain
@@
In this subcase, for ${\epsilon}=\pm 1$, by the definition in [1] and [2] the quadratic form $Q_{{\epsilon}}$ is given by $Q_{{\epsilon}}={\epsilon}(Q_{0}+p_{d}^{2}/2)$. Let ${\gamma}_{{\epsilon}}(t)$ be the fundamental solution of the initial value problem of the linear Hamiltonian system (3.16)-(3.17).
Then considering the above study on $Q_{0}$ on ${\bf R}^{2d}$ and $Q_{\pm}$ on ${\bf R}^{2}$, specially (3.3) and (3.10), by the study in Sections 1.8 and 1.9 on pp.36-47 in [15], the basic normal form decomposition of ${\gamma}_{\pm 1}$ can be written as follows,
@@
Because the invariants in (3.3) and the splitting numbers $S_{M}^{\pm}({\lambda})$ in [13] and Definition 9.1.4 on p.191 of [15] are symplectic additive, by (3.25) in the subcase 1.3.2 we obtain
@@
4 The (IR1) in Theorem 3.3 of [2] is a direct consequence of Remark 3.6 of [3]
In Section 3 of [2], as usual the authors denote the set of all non-degenerate symplectic matrices in ${\rm Sp}(2m)$ with $m\in{\bf N}$ by ${\rm Sp}^{\ast}(2m)$. Then for a symplectic path $\Phi$ with end point $\Phi(1)\in{\rm Sp}^{\ast}(2m)$, the end point matrix $\Phi(1)$ can be connected to I-I in ${\rm Sp}^{\ast}(2m)$ by a path $\Psi$ lying entirely in ${\rm Sp}^{\ast}(2m)$. Let $\Phi^{\prime}$ be the concatenating path of $\Phi$ and $\Psi$. Then as usual the Conley-Zehnder index of $\Phi$ is defined by
@@
Then the authors of [2] defined two indices for $\Phi\in\widetilde{{\rm Sp}}(2m)$ by
@@
Our next result shows that the relations of the property (IR1”) with the indices defined in [15] as well as [3] are as follows,
Proposition 2. (i) The assertion (IR1”-1) in (IR1) are direct consequence of (3.42)-(3.43) in [3].
(ii) The assertion (IR1”-2) in (IR1) was already proved in (1) of Theorem 1.1 of [8] (cf. $1^{0}$ of Theorem 10.1.2 on p.213 of [15]).
Proof. Note first that using definitions of $\mu_{\pm}(\Phi)$ in (4.2), by Corollary 6.1.12 and Definition 6.1.13 on p.144 of [15], it yields immediately